# An analytic redshift-independent formulation of baryonic effects on the matter power spectrum

### v1

@@ 1 Introduction
The Flamingo simulations have been shown to reproduce a series of observables of the galaxy and cluster population (Schaye et al., 2023; Braspenning et al., 2024; Braspenning et al., 2025; McCarthy et al., 2023, 2024; Kumar et al., 2025). As such, they are a great test-bed to measure the effect of baryons on the matter power spectrum. Furthermore, the use of variations of the base model, where the observables have been systematically shifted in proportion to the estimated errors on the galaxy stellar mass function and cluster gas fractions (Kugel et al., 2023), allows for a direct connection between the baryonic response – here in the form of $A_{\rm mod}(k)$ – and the observables the simulations were calibrated to. Additionally, the simulations themselves have already been used to investigate the role of baryons on the so-called S8 tension (McCarthy et al., 2023, 2024; Elbers et al., 2025; Schaller et al., 2025).
@@ 2 The FLAMINGO simulations
All the simulations used in this study were run in a volume of $1~{}\rm{Gpc}$ on the side, which is sufficient to obtain a converged baryon response of the matter power spectrum (Schaller et al., 2025). Apart from the runs used in Sec. 3.5, the simulations adopt as values of the cosmological parameters the maximum likelihood values from the DES year 3 data release (Abbott et al., 2022) combined with external probes [^1] Their “3$\times$2pt + All Ext.” model: $\Omega_{\rm m}=0.306,~{}\Omega_{\rm b}=0.0486,~{}\sigma_{8}=0.807,~{}h=0.681,~{}n_{\rm s}=0.967,\sum m_{\nu}c^{2}=~{}0.06{\rm eV}$.. The initial conditions (ICs) were generated using the MonofonIC code (Hahn et al., 2021; Elbers et al., 2022) using a 3-fluid formalism with a separate transfer function for each of the species.
@@ 3.3 The baryonic response
The fact that the response inferred from Flamingo deviates from the one obtained by Amon & Efstathiou (2022) implies that the simulations do not fit the data they used and do not resolve the S8 tension (See also McCarthy et al., 2023, 2024; Elbers et al., 2025). This is, at least in part, due to Amon & Efstathiou (2022) forcing the baryonic effects to solve the tension between KiDS lensing and the CMB and hence demanding a rather dramatic correction to the matter power spectrum whilst the Flamingo calibration strategy was to match the observed gas fraction in clusters. Note however, that the original tension between KiDS lensing and the CMB seems to have dissolved in more recent analysis of the weak-lensing data (Wright et al., 2025).
@@ 3.4 Evolution with redshift
As can be seen by comparing the solid lines in the different panels, the analytic correction displays only a small amount of evolution with redshift. The simulation results show a more significant evolution, especially at $z>1$. At $z\leq 1$, the analytic $A_{\rm mode}$ model matches the simulation results at the $2\%$ level up to scales $k=3~{}h\cdot{\rm Mpc}^{-1}$ for all the models shown here. We recall that the functional fit was only performed at $z=0$. It is thus quite remarkable that the simple one-parameter sigmoid is sufficient to capture the behaviour of simulations with different levels of baryonic response over a wide range of redshifts.
@@ 3.5 Background cosmology dependence
Figure 4: Same as Fig. 2 but for different background cosmologies in the Flamingo suite, as indicated in the legend. For the five cosmologies that do not involve decaying dark matter, the analytic formulation yields excellent results (the solid and dashed lines overlap). This confirms the assumption that the function $A_{\rm mod}(k)$ does not require a dependence on the choice of background cosmology to match our simulations. We caution, however, that the Flamingo suite only covers a small range of cosmology variations. See the text for the discussion of the two models including decaying dark matter.
@@
The fitting function we obtained above was constructed using Flamingo simulations that all adopt our fiducial cosmology (D3A). We now compare the fit to the other models that are part of the Flamingo suite. To this end, we keep $A_{\rm mod}$ as fitted in Sec. 3.2 but compute $P_{\rm m}^{\rm L}(k,z)$, $P_{\rm m}^{\rm NL}(k,z)$, and thus $\beta^{\rm NL}(k,z)$, from Class and the Mead et al. (2021) model for the different cosmologies in the Flamingo suite (See table 1 in Elbers et al., 2025, for the parameters of each model). We then obtain the baryonic response of the matter power spectrum and show the results using solid lines in Fig. 4. We compare our analytic model to the direct results of the simulations using dashed lines with the same colours.
Putting aside the two models featuring decaying dark matter (PlanckDCDM12 and PlanckDCDM24), we see that the analytic expression matches the results of the simulations extremely well. As we performed the fit only for our fiducial cosmology, this was not guaranteed a priori. We see also that there is very little dependence of the baryonic response on the cosmology. This was already noted by Schaller et al. (2025) (their Fig. 10) and can be explained by the model proposed by Elbers et al. (2025) who linked the changes in the response to changes in the halo mass-concentration relation. However, we caution that the range of cosmologies explored here does not include models that lead to large variations in this relation and thus in the response. We also note that we do not explore models that feature significant changes in the ratio $\Omega_{\rm b}/\Omega_{\rm m}$. We leave such explorations to future studies that make use of a wider range of cosmological models.
@@
Turning now our attention to the two models with decaying dark matter, we see that the analytic fit to $A_{\rm mod}$ predicts the same level of response as in the other cosmological models. This is in tension with the simulation results (dashed lines), which predict a rather different response. This is in part due to the non-linear model used to compute $P_{\rm m}^{\rm NL}$ not reproducing the simulations in this regime (see Appendix A) and in part due to the assumption, made when constructing $A_{\rm mod}$, that the correction does not depend on cosmology breaking down. This can be verified by using the results of the DMO model as input $P_{\rm m}^{\rm NL}$. When doing this, the analytic approximation also differs from the hydrodynamical simulation results (not shown).

### v2

@@ 1 Introduction
The Flamingo simulations have been shown to reproduce a series of observables of the galaxy and cluster population (Schaye et al., 2023; Braspenning et al., 2024; Braspenning et al., 2025; McCarthy et al., 2023, 2025; Kumar et al., 2025). As such, they are a great test-bed to measure the effect of baryons on the matter power spectrum. Furthermore, the use of variations of the base model, where the observables have been systematically shifted in proportion to the estimated errors on the galaxy stellar mass function and cluster gas fractions (Kugel et al., 2023), allows for a direct connection between the baryonic response – here in the form of $A_{\rm mod}(k)$ – and the observables the simulations were calibrated to. Additionally, the simulations themselves have already been used to investigate the role of baryons on the so-called S8 tension (McCarthy et al., 2023, 2025; Elbers et al., 2025; Schaller et al., 2025).
@@ 2 The FLAMINGO simulations
All the simulations used in this study were run in a volume of $1~{}\rm{Gpc}$ on the side, which is sufficient to obtain a converged baryon response of the matter power spectrum (Schaller et al., 2025). Apart from the runs used in Sec. 3.5, the simulations adopt as values of the cosmological parameters the maximum likelihood values from the DES year 3 data release (Abbott et al., 2022) combined with external probes, i.e. their “3$\times$2pt + All Ext.” model  [^1] The parameter values are given in the first row of Table 2..
The initial conditions (ICs) were generated using the MonofonIC code (Hahn et al., 2021; Elbers et al., 2022) using a 3-fluid formalism with a separate transfer function for each of the species.
@@ 3.3 The baryonic response
The fact that the response inferred from Flamingo deviates from the one obtained by Amon & Efstathiou (2022) implies that the simulations do not fit the data they used and do not resolve the S8 tension (See also McCarthy et al., 2023, 2025; Elbers et al., 2025). This is, at least in part, due to Amon & Efstathiou (2022) forcing the baryonic effects to solve the tension between KiDS lensing and the CMB and hence demanding a rather dramatic correction to the matter power spectrum whilst the Flamingo calibration strategy was to match the observed gas fraction in clusters. Note however, that the original tension between KiDS lensing and the CMB seems to have dissolved in more recent analysis of the weak-lensing data (Wright et al., 2025).
@@ 3.4 Evolution with redshift
As can be seen by comparing the solid lines in the different panels, the analytic correction displays only a small amount of evolution with redshift. The simulation results show a more significant evolution, especially at $z>1$. At $z\leq 1$, the analytic $A_{\rm mod}$ model matches the simulation results at the $2\%$ level up to scales $k=3~{}h\cdot{\rm Mpc}^{-1}$ for all the models shown here. We recall that the functional fit was only performed at $z=0$. It is thus quite remarkable that the simple one-parameter sigmoid is sufficient to capture the behaviour of simulations with different levels of baryonic response over a wide range of redshifts.
@@ 3.5 Background cosmology dependence
Table 2: The values of the cosmological parameters for the different simulations of the Flamingo suite used in this study. All simulations assume a flat $\Lambda$CDM Universe including massive neutrinos with $N_{\rm eff}=3.044$ effective relativistic neutrino species at high redshift and with an amount of radiation corresponding to $T_{\rm CMB}=2.7255~{}{\rm K}$. For the simulations with decaying cold dark matter (DCDM), the $\Omega_{\rm cdm}$ corresponds to the sum of the present-day densities of decaying cold dark matter and dark radiation. For these models, the last column shows the dark matter decay rate, $\Gamma$, in units of $100~{}{\rm km/s/Mpc}=H_{0}/h$.
Simulation name $h$ $\Omega_{\text{m}}$ $\Omega_{\text{cdm}}$ $\Omega_{\text{b}}$ $\sum m_{\nu}$ $\sigma_{8}$ $10^{9}A_{\text{s}}$ $n_{\text{s}}$ $\Gamma h/H_{0}$
[…]
Figure 4: Same as Fig. 2 but for different background cosmologies in the Flamingo suite, as indicated in the legend both at $z=0$ (left) and $z=1$ (right). For the five cosmologies that do not involve decaying dark matter, the analytic formulation yields excellent results (the solid and dashed lines overlap), especially at $z=0$. This confirms the assumption that the function $A_{\rm mod}(k)$ does not require a dependence on the choice of background cosmology to match our simulations. We caution, however, that the Flamingo suite only covers a small range of cosmology variations. See the text for the discussion of the two models including decaying dark matter.
@@
The fitting function we obtained above was constructed using Flamingo simulations that all adopt our fiducial cosmology (D3A). We now compare the fit to the other models that are part of the Flamingo suite. To this end, we keep $A_{\rm mod}$ as fitted in Sec. 3.2 but compute $P_{\rm m}^{\rm L}(k,z)$, $P_{\rm m}^{\rm NL}(k,z)$, and thus $\beta^{\rm NL}(k,z)$, from Class and the Mead et al. (2021) model for the different cosmologies in the Flamingo suite whose parameter values are given in Table 2. We then obtain the baryonic response of the matter power spectrum and show the results at $z=0$ using solid lines in the left panel of Fig. 4. We compare our analytic model to the direct results of the simulations using dashed lines with the same colours.
Putting aside the two models featuring decaying dark matter (PlanckDCDM12 and PlanckDCDM24), we see that the analytic expression matches the results of the simulations extremely well. As we performed the fit only for our fiducial cosmology, this was not guaranteed a priori. We see also that there is very little dependence of the baryonic response on the cosmology. This was already noted by Schaller et al. (2025) (their Fig. 10) and can be explained by the model proposed by Elbers et al. (2025) who linked the changes in the response to changes in the halo mass-concentration relation. However, we caution that the range of cosmologies explored here does not include models that lead to large variations in this relation and thus in the response. Two important type of parameter variations have nevertheless been explored: those leading to scale-dependent (changes in $n_{\rm s},\sum m_{\nu}$) and scale-independent (changes in $\Omega_{\rm m},\sigma_{8}$) effects on the linear power spectrum.
We also note that we do not explore models that feature significant changes in the ratio $\Omega_{\rm b}/\Omega_{\rm m}$. We leave such explorations to future studies that make use of a wider range of cosmological models.
@@
On the scale of interest, $\beta^{\rm NL}$ is dominated by the one-halo term, which displays little cosmology-dependence in halo models. It is thus interesting to repeat the comparison at higher redshift where this term is smaller. We show the response obtained in the simulations and using our analytic model at $z=1$ in the right panel of Fig. 4. For the simulations (dashed lines), a greater level of cosmology-dependence than at $z=0$ can be seen. This is, however, not captured by the analytic model (solid lines). The halo model does not capture the cosmology-dependence of the one-halo to two-halo transition. More advanced modeling of $\beta^{\rm NL}$ in the future might reconcile the simulations and the model. We note, however, that the differences seen here are nonetheless only at the $2\%$ level for $k<3~{}h\cdot{\rm Mpc}^{-1}$.
Turning now our attention to the two models with decaying dark matter, we see that the analytic fit to $A_{\rm mod}$ predicts the same level of response as in the other cosmological models. This is in tension with the simulation results (dashed lines), which predict a rather different response at $z=0$. This is in part due to the non-linear model used to compute $P_{\rm m}^{\rm NL}$ not reproducing the simulations in this regime (see Appendix A) and in part due to the assumption, made when constructing $A_{\rm mod}$, that the correction does not depend on cosmology breaking down. This can be verified by using the results of the DMO model as input $P_{\rm m}^{\rm NL}$. When doing this, the analytic approximation also differs from the hydrodynamical simulation results (not shown). At $z=1$ (right panel), the responses extracted from the simulations with decaying dark matter display a behaviour closer to the other cosmologies, and is hence captured better by our analytic model than at $z=0$.
@@ Acknowledgements
We thank the anonymous referee for their helpful comments.